# Hierarchy-Aware Multimodal Unlearning for Medical AI

### v2

@@ 1 Introduction
We evaluate hierarchy-aware multimodal unlearning across four levels of medical structure and across three tasks in MedForget: generation, cloze completion and classification evaluated using generation quality, cloze reasoning, and classification accuracy respectively. Results reveal a consistent deletion–utility trade-off among existing methods: training-based approaches preserve retain and general medical performance but leave substantial residual memorization on the forget set, while aggressive training-free methods achieve stronger deletion at the cost of significant utility degradation. In contrast, CHIP consistently achieves the largest forget–retain performance gap across hierarchies, yielding the lowest forget-set scores while maintaining competitive retain performance. Even when a specific item is unlearned, the remaining correlated context can enable implicit reconstruction of the deleted information. Flat benchmarks fail to expose such leakage pathways and therefore overestimate unlearning effectiveness, highlighting the need for structure-aware multimodal unlearning evaluation frameworks and methods, particularly in regulated domains such as healthcare under HIPAA (Act, 1996) and GDPR (GDPR, 2016). Our results show that CHIP is substantially more robust to such hierarchical reconstruction attacks, exhibiting lower leakage even when adversaries exploit rich contextual identifiers. Across tasks, CHIP also demonstrates more uniform forgetting under comparable generation quality constraints, particularly for cloze and classification tasks. Together, these results show that hierarchy-aware unlearning is both necessary and challenging in medical MLLMs, and that CHIP provides a more effective and robust balance between privacy protection and downstream utility than existing approaches.
@@ 2 Related Work
Machine unlearning aims to remove the influence of specific data subsets from trained models without retraining from scratch. Existing unlearning benchmarks primarily focus on unimodal data (Thudi et al., 2022; Patil et al., 2024a, 2025), leaving multimodal unlearning largely unexplored. Text-based datasets such as TOFU (Maini et al., 2024) and MUSE (Shi et al., 2023) evaluate selective forgetting in LLMs but do not capture cross-modal dependencies. There are works (Choi et al., 2024; Dahal and Xiong, 2025) showing that existing LLM unlearning methods fail to reliably remove structured, multi-hop knowledge, whether through indirect reasoning chains or knowledge graph relations, highlighting that flat unlearning is insufficient. A few recent efforts explore multimodal unlearning (Liu et al., 2025a; Patil et al., 2024b), but they primarily evaluate flat unlearning without modeling the hierarchical relationships between data sources or tasks, which fail to reflect how deletions naturally arise in clinical workflows.
@@
Regulations (Act, 1996; GDPR, 2016) mandate strict control over identifiable health information and support the "right to be forgotten," motivating techniques that can selectively remove private data from trained models. Although large-scale datasets like MIMIC-CXR (Johnson et al., 2019) have catalyzed progress in multimodal learning for radiology (Hartsock and Rasool, 2024; Li et al., 2023), they also expose risks of patient re-identification and data leakage through inference or inversion attacks (Shokri et al., 2017; Jagielski et al., 2020). Prior defenses such as anonymization, differential privacy, and federated learning (McMahan et al., 2017) mitigate but do not eliminate these risks. Recent research in machine unlearning (Xu et al., 2023; Wang et al., 2023; Zhang et al., 2023b; Campagnano, 2024; Liu et al., 2025a) offers a more targeted solution by enabling models to forget specific data without retraining. Existing datasets overlook the multimodal, hierarchical structure of medical data, where forgetting at one level induces cascading effects. We address this with a hierarchy-aware unlearning framework.
@@ 3 Benchmark: MedForget
In this section, we first describe the hierarchical structure and unlearning design in  Section˜3.1.  Section˜3.2 then details the dataset construction, including data collection and task design.  Section˜3.3 explains the dataset partitioning strategy. Finally, Section˜3.4 presents key dataset statistics.
@@ 3.3 Dataset Partition
Figure 5: Illustration of CHIP. The method (a) collects cross-modal activations, (b) identifies target-related neurons, (c) constructs hierarchy-aware directions that isolate sibling-differential information, and (d) applies a training-free subspace projection to remove these directions from both languages and vision–language weights.
@@ 3.4 Dataset Statistics
MedForget contains 8 institutions, 64 patients, 256 studies, and 1,024 report sections, yielding 3,840 multimodal VQA pairs. We apply a 25% forget ratio at each hierarchy level, resulting in 960 forget and 2,880 retain VQA pairs. The dataset includes 3,072 section-level tasks (generation, cloze, classification) and 768 hierarchical identity tasks evenly split across institution, patient, and study levels. At the section level, the dataset reflects radiology report composition: Impression (23.7%), Findings (20.1%), Technique (19.6%), Indication (17.1%), and Examination (13.4%) are most frequent, while clinically rare sections (e.g., Clinical History, Notification) occur infrequently (full distribution shown in Figure˜4). See Section˜A.10 for detailed dataset statistics.
@@ 4 Method: CHIP
Not all neurons contribute equally to storing forget-specific information. Inspired by Liu (2025), we identify neurons most relevant to forgetting via importance scores. Let $\mathbf{I}_{\text{forget}}^{(l)}$ and $\mathbf{I}_{\text{retain}}^{(l)}$ denote mean absolute activations over forget and retain samples at layer $l$. We compute $s_{i}^{(l)}=\mathbf{I}_{\text{forget},i}^{(l)}-\mathbf{I}_{\text{retain},i}^{(l)}$ and select the top $k\%$ neurons with the highest scores as set $\mathcal{S}^{(l)}$. These selected neurons form the subspace where surgery will be applied (see Figure˜5 (b)).
@@ 5.1 Experimental Setup
We evaluate unlearning performance using three complementary metrics that capture both forgetting completeness and retain set utility: (1) Generation Score (Gen Score) for generation tasks, computed as a weighted average of 75% factuality score and 25% ROUGE-L (Lin, 2004). The factuality score is obtained by using DeepSeek-V3 (DeepSeek-AI et al., 2025) as LLM-as-a-judge (Gu et al., ) to rate clinical accuracy on a 1–10 scale, following prior work (Sun et al., 2023; Yu et al., 2024; Zheng et al., 2023); (2) Cloze Accuracy (Cloze Acc), measured by exact string match in cloze-style completion tasks; and (3) Classification Accuracy (Class. Acc), computed as the proportion of correct multiple-choice predictions given the question and chest X-ray image. Evaluation metric details are provided in Appendix A.3.
@@ Appendix A Appendix
10–9:
Fully correct; all key findings present
Table 4: Prompt template for factuality evaluation.
@@ A.1 Compared Unlearning Methods
  • Gradient Difference (Yao et al., 2024): an extension of gradient ascent that jointly optimizes forgetting and retention objectives. Unlike vanilla gradient ascent, which only increases the loss on the forget set, Grad-Diff introduces a balanced objective that simultaneously decreases loss on the retain set. The objective is given by
@@
    where $S_{F}$ and $S_{R}$ denote the forget and retain sets, respectively. During training, examples from both sets are interleaved to maintain computational efficiency while ensuring that unlearning does not degrade model utility on retained data.
  • KL Minimization (Nguyen et al., 2020): a divergence-guided unlearning technique that promotes targeted forgetting on the forget set $S_{F}$ while anchoring the model’s outputs on the retain set $S_{R}$ to those of the original fine-tuned model. This approach maximizes the cross-entropy loss on $S_{F}$ to disrupt memorization and minimizes the Kullback-Leibler (KL) divergence on $S_{R}$ to safeguard retained knowledge. The composite objective is given by
@@
  • Negative Preference Optimization (NPO) (Zhang et al., 2024): a preference-inspired optimization framework designed to mitigate instability in direct loss maximization on forget data. Drawing from direct preference optimization principles (Rafailov et al., 2024), NPO treats forget set samples as negative preferences, encouraging the model to downweight their influence relative to a reference policy derived from retain set fine-tuning. The loss is formulated as
@@
  • Modality-Aware Neuron Unlearning (MANU) (Liu et al., 2025b): a neuron-level unlearning framework for multimodal large language models (MLLMs) that selectively removes modality-specific knowledge contributions. MANU operates in two stages: (1) important neuron selection, which computes an aggregated importance score $\mathcal{I}(\mathcal{D},n)$ for each neuron $n$ across modalities using four metrics—absolute ($I_{\text{abs}}$), frequency ($I_{\text{freq}}$), variance ($I_{\text{var}}$), and root mean square ($I_{\text{rms}}$)—defined on activation deviations between textual and multimodal subsets of $\mathcal{D}$; and (2) selective pruning, which ranks neurons by the ratio score $S_{n}=\mathcal{I}(\mathcal{D}_{f},n)/(\mathcal{I}(\mathcal{D}_{r},n)+\epsilon)$ and sets weights to zero for the top $\alpha\%$ most forget-associated neurons. By disentangling modality-specific activations, MANU enables fine-grained forgetting across both textual and visual streams, achieving balanced unlearning with minimal performance degradation on retain data.
@@ A.5 Qualitative Examples
Figure 9: More examples of data samples in MedForget.
@@ A.7.3 Weight Modification
We compare our subspace projection approach against the zero-out pruning strategy. While both methods modify weights based on importance scores, projection removes only the forget-specific subspace components, whereas zero-out completely eliminates selected neuron contributions.
@@
(3) The performance gap is substantial across all granularities. The gap between CHIP and zero-out ranges from 10.63 points at Section level to 13.65 points at Institution level. This consistent advantage demonstrates that subspace projection is fundamentally more suitable for selective unlearning than zero-out pruning, regardless of the hierarchy granularity.
A.8 Broader Impact
MedForget addresses critical privacy challenges in medical AI by introducing the first comprehensive benchmark for hierarchy-aware unlearning in multimodal clinical models. By explicitly modeling realistic medical hierarchies (e.g., institution, patient, study, and section), our benchmark supports the development and evaluation of unlearning methods aligned with regulatory requirements such as HIPAA (Act, 1996) and GDPR (GDPR, 2016), including the “right to be forgotten.” Our findings expose fundamental trade-offs between deletion granularity, forgetting completeness, and retained clinical utility, providing actionable insights for deploying privacy-preserving medical MLLMs in practice. The proposed hierarchy-aware unlearning approach further enables more selective removal of sensitive information while minimizing collateral damage to retained data and downstream performance. At the same time, our results highlight that unlearning is not a complete substitute for broader privacy protections: fine-grained deletion remains vulnerable to contextual reconstruction attacks, and unlearning guarantees depend on model architecture and threat assumptions. We therefore view this work as a step toward safer medical AI systems, complementing—rather than replacing—other safeguards such as access control, auditability, and data minimization.
A.9 Potential Risks
This work has potential risks that should be considered when deploying hierarchy-aware unlearning in medical MLLMs. First, incomplete or incorrect hierarchical metadata may lead to insufficient forgetting or unintended deletion of non-target information, potentially violating privacy guarantees or harming clinical utility. Since real-world medical hierarchies are often noisy or inconsistently defined, errors in structure could propagate into unlearning decisions.
Second, although our approach reduces leakage under hierarchical reconstruction attacks, it does not fully eliminate the risk of adversarial recovery using auxiliary knowledge, external databases, or unseen correlations beyond the modeled hierarchy. Such attacks may still enable partial reconstruction of sensitive medical content.
Finally, selective unlearning mechanisms may create a false sense of compliance if used without rigorous auditing. Claims of successful deletion should be validated with diverse evaluation protocols, as residual memorization may persist despite favorable aggregate metrics. Careful monitoring and verification are therefore critical when applying unlearning methods in high-stakes domains such as healthcare.
A.10 Dataset Statistics
@@
A.11 LLM Usage
We used large language models as writing assistants for grammar correction, clarity improvement, and minor rewriting of text. All technical content, experimental design, analysis, and conclusions were developed by the authors.

### v3

@@ 1 Introduction
We evaluate hierarchy-aware multimodal unlearning across four levels of medical structure and across three tasks in MedForget: generation, cloze completion and classification evaluated using generation quality, cloze reasoning, and classification accuracy respectively. Results reveal a consistent deletion–utility trade-off among existing methods: training-based approaches preserve retain and general medical performance but leave substantial residual memorization on the forget set, while aggressive training-free methods achieve stronger deletion at the cost of significant utility degradation. In contrast, CHIP consistently achieves the largest forget–retain performance gap across hierarchies, yielding the lowest forget-set scores while maintaining competitive retain performance. Even when a specific item is unlearned, the remaining correlated context can enable implicit reconstruction of the deleted information. Flat benchmarks fail to expose such leakage pathways and therefore overestimate unlearning effectiveness, highlighting the need for structure-aware multimodal unlearning evaluation frameworks and methods, particularly in regulated domains such as healthcare under HIPAA (Act, 1996) and GDPR (Regulation, 2016). Our results show that CHIP is substantially more robust to such hierarchical reconstruction attacks, exhibiting lower leakage even when adversaries exploit rich contextual identifiers. Across tasks, CHIP also demonstrates more uniform forgetting under comparable generation quality constraints, particularly for cloze and classification tasks. Together, these results show that hierarchy-aware unlearning is both necessary and challenging in medical MLLMs, and that CHIP provides a more effective and robust balance between privacy protection and downstream utility than existing approaches.
@@ 2 Related Work
Machine unlearning aims to remove the influence of specific data subsets from trained models without retraining from scratch. Existing unlearning benchmarks primarily focus on unimodal data (Thudi et al., 2022; Patil et al., 2024a, 2025), leaving multimodal unlearning largely unexplored. Text-based datasets such as TOFU (Maini et al., 2024) and MUSE (Shi et al., 2025) evaluate selective forgetting in LLMs but do not capture cross-modal dependencies. There are works (Choi et al., 2024; Dahal and Xiong, 2025) showing that existing LLM unlearning methods fail to reliably remove structured, multi-hop knowledge, whether through indirect reasoning chains or knowledge graph relations, highlighting that flat unlearning is insufficient. A few recent efforts explore multimodal unlearning (Liu et al., 2025a; Patil et al., 2024b), but they primarily evaluate flat unlearning without modeling the hierarchical relationships between data sources or tasks, which fail to reflect how deletions naturally arise in clinical workflows.
@@
Regulations (Act, 1996; Regulation, 2016) mandate strict control over identifiable health information and support the "right to be forgotten," motivating techniques that can selectively remove private data from trained models. Although large-scale datasets like MIMIC-CXR (Johnson et al., 2019) have catalyzed progress in multimodal learning for radiology (Hartsock and Rasool, 2024; Li et al., 2023), they also expose risks of patient re-identification and data leakage through inference or inversion attacks (Shokri et al., 2017; Jagielski et al., 2020). Prior defenses such as anonymization, differential privacy, and federated learning (McMahan et al., 2017) mitigate but do not eliminate these risks. Recent research in machine unlearning (Nguyen et al., 2025; Jang et al., 2023; Zhang et al., 2023b; Trippa et al., 2024; Liu et al., 2025a) offers a more targeted solution by enabling models to forget specific data without retraining. Existing datasets overlook the multimodal, hierarchical structure of medical data, where forgetting at one level induces cascading effects. We address this with a hierarchy-aware unlearning framework.
@@ 3 Benchmark: MedForget
Here, we describe the hierarchical structure and unlearning design in  Section˜3.1.  Section˜3.2 then we talk about the dataset construction, including data collection and task design. In  Section˜3.3, we explain the dataset partitioning strategy. Finally, we present key dataset statistics in Section˜3.4.
@@ 3.3 Dataset Partition
Figure 5: Illustration of CHIP. The method (a) collects cross-modal activations, (b) identifies target-related neurons, (c) constructs hierarchy-aware directions that isolate sibling-differential information, and (d) applies a training-free subspace projection to remove these directions from both language and vision–language weights.
@@ 3.4 Dataset Statistics
MedForget contains 8 institutions, 64 patients, 256 studies, and 1,024 report sections, yielding 3,840 multimodal VQA pairs. We apply a 25% forget ratio at each hierarchy level, resulting in 960 forget and 2,880 retain VQA pairs. The dataset includes 3,072 section-level tasks (generation, cloze, classification) and 768 hierarchical identity tasks evenly split across institution, patient, and study levels. At the section level, the dataset reflects radiology report composition: Impression (23.7%), Findings (20.1%), Technique (19.6%), Indication (17.1%), and Examination (13.4%) are most frequent, while clinically rare sections (e.g., Clinical History, Notification) occur infrequently (full distribution shown in Figure˜4). See Section˜A.8 for detailed dataset statistics.
@@ 4 Method: CHIP
Not all neurons contribute equally to storing forget-specific information. Inspired by Liu et al. (2025b), we identify neurons most relevant to forgetting via importance scores. Let $\mathbf{I}_{\text{forget}}^{(l)}$ and $\mathbf{I}_{\text{retain}}^{(l)}$ denote mean absolute activations over forget and retain samples at layer $l$. We compute $s_{i}^{(l)}=\mathbf{I}_{\text{forget},i}^{(l)}-\mathbf{I}_{\text{retain},i}^{(l)}$ and select the top $k\%$ neurons with the highest scores as set $\mathcal{S}^{(l)}$. These selected neurons form the subspace where surgery will be applied (see Figure˜5 (b)).
@@ 5.1 Experimental Setup
We evaluate unlearning performance using three complementary metrics that capture both forgetting completeness and retain set utility: (1) Generation Score (Gen Score) for generation tasks, computed as a weighted average of 75% factuality score and 25% ROUGE-L (Lin, 2004). The factuality score is obtained by using DeepSeek-V3 (DeepSeek-AI et al., 2025) as LLM-as-a-judge (Gu et al., 2024) to rate clinical accuracy on a 1–10 scale, following prior work (Sun et al., 2023; Yu et al., 2024; Zheng et al., 2023); (2) Cloze Accuracy (Cloze Acc), measured by exact string match in cloze-style completion tasks; and (3) Classification Accuracy (Class. Acc), computed as the proportion of correct multiple-choice predictions given the question and chest X-ray image. Evaluation metric details are provided in Appendix A.3.
@@ A.1 Compared Unlearning Methods
Gradient Difference (Yao et al., 2024): It is a version of gradient ascent that optimizes the forget and retain objectives. Grad-Diff introduces a balanced objective that simultaneously decreases loss on the retain set. The objective is given by
@@
where $S_{F}$ and $S_{R}$ denote the forget and retain sets, respectively. We interleave samples from both sets for computational efficiency while ensuring that deletion does not damage model utility on the retain data.
KL Minimization (Nguyen et al., 2020): It an unlearning technique that is guided by divergence and aims for targeted forgetting on the forget set $S_{F}$. It simultaneously aims to anchor the model’s outputs on the retain set $S_{R}$ to those of the original fine-tuned model. Its approach aims to maximize the cross-entropy loss on $S_{F}$ for deletion and aims to minimize theKL divergence on $S_{R}$ to get high retain set performance. The composite objective is given by
@@
Negative Preference Optimization (NPO) (Zhang et al., 2024): It is a preference optimization method designed to reduce the instability in gradient ascent. NPO treats forget set samples as negative preferences, encouraging the model to downweigh their influence relative to a reference policy derived from the retain set. The loss is
@@
Modality-Aware Neuron Unlearning (MANU) (Liu et al., 2025b): It is a neuron-level pruning for unlearning framework for MLLMs that selectively removes modality-specific knowledge contributions. It works in two stages: (1) important neuron selection, where it computes an importance score $\mathcal{I}(\mathcal{D},n)$ for each neuron $n$ across both modalities using four metrics, absolute ($I_{\text{abs}}$), frequency ($I_{\text{freq}}$), variance ($I_{\text{var}}$), and root mean square ($I_{\text{rms}}$), defined on activation deviations between textual and multimodal subsets of $\mathcal{D}$; and (2) selective pruning, which ranks neurons by the ratio score $S_{n}=\mathcal{I}(\mathcal{D}_{f},n)/(\mathcal{I}(\mathcal{D}_{r},n)+\epsilon)$ and sets weights to zero for the top $\alpha\%$ most forget-associated neurons. By disentangling modality-specific activations, it aims to enable targeted forgetting across both modalities, achieving balanced unlearning.
@@ A.3 Evaluation Metrics
Figure 9: More examples of data samples in MedForget.
@@ A.7.3 Weight Modification
We compare our subspace projection approach against the zero-out pruning strategy. While both methods modify weights based on importance scores, projection removes only the forget-specific subspace components, whereas zero-out eliminates selected neuron contributions.
@@
(3) The performance gap is substantial across all granularities. The gap between CHIP and zero-out ranges from 10.63 pts at Section level to 13.65 pts at Institution level. This consistent advantage shows that subspace projection is fundamentally more suitable for selective unlearning than zero-out pruning, regardless of the hierarchy granularity.
A.8 Dataset Statistics
@@
10–9:
Fully correct; all key findings present
Table 7: Prompt template for factuality evaluation.